OpenStudy (anonymous):

Taylor is making a mini apple pie with a radius of 3.14 in for her math class. What is the area of her mini pie? A. 31 in^2 B. 32 in^2 C. 30 in^2 D. 15 in^2

OpenStudy (anonymous):

A = (pi)(r^2) In this problem, your r is "pi", so you have: A = (pi)[(pi)^2] = (pi)^3 You can use 3.14 for pi and round the answer.

OpenStudy (anonymous):

I got A as my answer

OpenStudy (anonymous):

is that correct @tcarroll010 ?

OpenStudy (anonymous):

Yes, you are right!

OpenStudy (anonymous):

thx for the recognition!

OpenStudy (anonymous):

No problem. I would have had the same answer. You bat me to it though. Haha